OpenStudy (anonymous):

27 equal cubes, each with a side length r, are arranged so as to form a single larger cube with a volume of 81. If the larger cube has a side length of s, then what is r divided by s?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

how did u do that

OpenStudy (anonymous):

s^3 =81 then s=3(3)^1/3 its possible when r=3^1/3

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

yeahthanksalot